for certain organ pipe three successive resonance frequency are 425,595,765Hz.if speed of air iw 340.find length of pipe?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Dear Parul,
Already answered this few minutes before,
plz check your previous sn.
But for ease I am posting the naswer again:
 
Assuming to be open pipe the Fundamental frequency needs to be calculated first as:
For open organ pipe,
the frequency ratio is: 1:3:5..............
hence,
595/425 = 7/5
ANd therfore, the fundamental frequence would be = > 425/5 = 85.
85 = v/2l
2l = v/85=> 340/85
l  = 340/(85*2) = 2 meter.